# arrows,decorations.markings,shapes,matrix,positioning, calc, fit,decorations.pathreplacing,angles,quotes,backgrounds

### v1

@@ 7.1. Justification Involving Protected Feature
Here, we analyze justifications that explicitly mention “gender”, “male”, “female”, etc. in the direct bias conditions. First, we consider such justifications before we disclose model bias to participants. In this phase, when we do not provide participants with explanations, we almost never observe participants directly discussing “gender” as a salient part of their justification. When they are given explanations, participants mention trying to “ignore gender” when making their decision. We observe that participants who mentioned gender in a prediction about a female applicant tended to be making “Complete” or “Neutral” predictions. Thus, even though explanations significantly decreased gender parity (§6.1) overall, they do seem to help certain participants correct model biases in some cases. After bias disclosure, justifications mentioning gender still predominantly appear in the condition with explanations. However, we do find some participants who, even without explanations, mention gender bias and flip “Late” predictions for female applicants. For example, one participant explained they overrode such a prediction because they had seen an example of a male applicant with the same occupation who was predicted as “Complete”. In a condition with explanations, many participants mention that “gender should not be a deciding factor” and that they try to ignore this feature when making their prediction. Some participants even explain when they are agreeing with a “Late” prediction for a female applicant that their decision was based on other features (“The large amount of negatives aside from gender still point towards being late.”). However, even in the case of direct bias and with both explanations and bias disclosure, we still observe participants who agree with model biases. For example, one participant agreed with a “Complete” prediction of a male applicant “[b]ecause according to AI, male gender is more likely to complete loan…
@@ 7.2. Justifications Involving Proxy Feature
Here, we analyze justifications that explicitly mention “university”, “college”, “school”, etc. in the indirect bias conditions. First, we consider how participants used the “university” feature in their justifications before any disclosures of model bias or correlation. Overall, some participants mention that attending college in the first place generally increases the likelihood of repayment, regardless of the specific school (for example, mentioning that an applicant is “college educated” and predicting “Complete”). However, other participants do consider the specific college as a reason to accept or reject an applicant. In the conditions without explanations, we see evidence of participants relying on their own judgment of school quality. In the underlying model, going to a co-ed school is counted as a positive. We instead see participants mentioning that they have “Never heard of Kenyon College,” or that the co-ed school in the application “…is not a particularly prestigious university.” In contrast, when participants are provided explanations, we instead see examples of participant’s evaluation of a school aligning with the model’s biases. For example, on applicants from women’s colleges, participants claimed “The applicant didn’t go to a good college,” or that “…College history was a major contributing factor to being late on loan,” while on applications from co-ed schools, participants claimed that the applicant “…attended a good university,” or mention that “the university is listed as a good one.” This supports our quantitative finding that explanations alone lead participants to be more influenced by model biases in their decision-making (§6.1). Next, we consider how participants talk about the “university” feature after disclosures. With bias disclosure alone (and especially without explanations), mentions of university were quite sparse. Some participants mention that university is given too much weight (“I just find it hilarious that the borrower’s state and university is such a huge factor.”) but may not make the connection that this over-use is due to indirect bias. One participant knew outside of the study that Bryn Mawr is a women’s college, but they were still generally not confident about how to find biased predictions without direct access to protected features: “After learning more about possible discriminatory predictions on the AI’s part… I’m specifically concerned about gender and race… but don’t quite know how to discern that from these charts… This applicant profile gave me pause because I *think* Bryn Mawr College is an all-women’s college. When participants were instead given the full bias and correlation disclosure, the “university” feature appears frequently in decision justifications. Here, participants continue to point out that university is given too much weight in explanations generally but also point to undue negative weight towards women’s colleges. (e.g., “While the system said late, I thought this was unfair because it placed a strong negative value on the college, which might be a women’s college.”). While some participants bring up this university bias as a justification for flipping model predictions, many acknowledge the bias and make a neutral prediction or agree with predictions of women being late in repayment. We also find that some participants struggled to recall which universities were co-ed, and others misinterpreted the co-ed schools as being predominantly male which may have limited their ability to intervene to correct model biases.
@@ 7.3. Justifications Indicating Over-reliance
In addition to the positive examples of explanations and disclosures helping participants notice and correct model biases, in many instances, as gleaned through their justifications, we also find instances of participants making predictions entirely based on the AI prediction or the corresponding explanations, regardless of disclosures. For instance, even after bias disclosure, one participant in a direct bias condition agreed with a prediction of a female applicant being “Late” saying that “They seem to have more negatives than positives.” Similarly, even after bias and correlation disclosures, a participant in an indirect bias condition changed the prediction for an applicant from a women’s college from “Complete” to “Late”, making a similar justification. For the applicants in question, however, if the biased features (gender and university, respectively) were to be ignored, the positives in fact would outweigh the negatives. This indicates that some participants continue to focus their decision-making process on explanations that they know contain biases. We also find instances of over-trust in AI even after being told that AI is biased such as “I have no reason to disagree with the AI, if the AI is discriminating it probably has a good reason to,” or “An AI is usually better than a professional let alone an amateur like me.” This indicates that even despite explanations and disclosures, there is much room for improvement in educating and training humans to avoid unwarranted trust in AI systems and promote fair decision-making.
@@ 8. Discussion and Limitations
Explanations Only Disclosures Without Explanations Disclosures With Explanations Joint Intervention Prot Prox Prot BD Prox BD Prox BD+CD Prot BD Prox BD Prox BD+CD Prot BD Prox BD Prox BD+CD Fairness Rating $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ Fairness Perception Fairness Saliency $\uparrow$ $\cdot$ $\cdot$ $\cdot$ $\uparrow$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\cdot$ $\uparrow$ DM Fairness Gender Parity $\downarrow$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\cdot$ Accuracy $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\downarrow$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$ FNR $\uparrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ DM Quality FPR $\downarrow$ $\uparrow$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\cdot$ Table 1. Summary of our main results. Arrows represent significant effects and point in the direction of the change. “BD” and “CD” represent bias and correlation disclosures, respectively. In this work, we studied the effect of explanations and disclosures on fairness perceptions and decision-making when humans are provided predictions from models exhibiting direct or indirect bias. Our findings are summarized in Table 1. Regardless of intervention, we consistently observed that human-AI teams made fairer decisions than the AI alone. We found that explanations alone significantly improved participants’ ability to notice unfairness in the case of direct bias only. However, explanations led participants to be more influenced by model biases, whether they noticed these biases or not. Disclosures were an effective tool for helping users recognize unfairness in the case of indirect bias, especially with the help of explanations. And we saw that this increased recognition of bias was paired with fairer human-AI decisions, showing that disclosures helped participants understand when and how to intervene on model decisions to produce fairer outcomes. However, we found that the joint intervention of including both explanations and disclosures (over including neither) was only effective in helping participants recognize model bias, not correct it. If the main objective is to help users notice model unfairness in the case of direct bias, we recommend including explanations, and if it is to help users notice model unfairness in the case of indirect bias, we recommend including explanations and disclosing both model bias and the correlations between protected and proxy features. But if the main objective is to help the human-AI team produce fairer outcomes, we did not find including explanations with disclosures to be an effective intervention. However, if explanations are to be used, then disclosures may help contextualize explanations and the potential biases, especially when these biases are indirect. While in a perfect world, such known biases could be addressed in the model itself instead of relying on human intervention, this may not always be possible. In many cases, we may have limited access to the underlying model (e.g., only having API access) or may not be able to non-superficially “debias” it (Gonen and Goldberg, 2019). Disclosures may help uncover these biases to humans, possibly leading to fairer human-AI decisions. A key limitation of work is that since we show our participants partially-synthetic loan data, we cannot directly rely on the existing ground truth. Instead, we calculate the expectation of ground-truth based metrics (accuracy, FNR, and FPR) which means that there are applicants for which neither choice is very likely to be “correct” (i.e., both $P(Y_{i}=1)$ and $P(Y_{i}=0)$ are close to $0.5$). We handle this in part by adjusting for the baseline AI-only scores; however, using a fully non-synthetic dataset and original ground truth values may lead to cleaner results. This lack of a true ground-truth, in part, led us to use demographic parity which has been argued to be insufficient as a notion of fairness (Dwork et al., 2012). Another limitation is that our study design forces participants to make decisions one at a time, without seeing the entire pool of applicants. It is our hope that the percent/percentile information given for each feature helped give participants a better sense of how each applicant’s profile compared to the general pool, even without seeing many profiles. However, we recognize that it may be difficult for participants to conceptualize what a “strong” or “weak” candidate looks like under this design. This may make it more difficult for participants who, for example, wish to increase the acceptance rate of women in phase 2 to decide which female applicants are “most deserving” of having their prediction flipped to “Complete”. Despite these limitations, our work provides insights into the effect of explanations on fairness in human-AI decision-making, especially when the biases are indirect (through a proxy features). We conclude that neither explanations nor disclosures alone improve the fairness of decisions made by a human-AI team. Our findings serve to caution the wider community from treating explanations as a foolproof solution to human-AI collaborative decision-making: explanations may not always make model biases clear and may make people more prone to align with model biases, leading to less fair decisions. When people are repeatedly exposed to explanations that justify or rationalize biased predictions, they may begin to accept these biases as valid or even desirable, rather than critically questioning and challenging them. We highlight that explanations and disclosures in conjunction may be helpful to some extent. However, more work is needed to further examine how best to aid humans not only in identifying indirect model biases, but also in systematically correcting these biases.
@@ Acknowledgements.
We sincerely thank the CHI TRAIT reviewers, Md Naimul Hoque, and the members of the UMD CLIP and HCIL labs for their valuable feedback.
@@ A.1. Probability of Loan Completion
In our study, each applicant $i$ is only evaluated by a single participant $j$ in a given condition, and each participant $j$ evaluates $20$ applicants across the two phases (§4.2). We represent the set of observed decisions as $S=\{(i,j)\mid\text{participant $j$ sees applicant $i$}\}$. For the $i^{\text{th}}$ applicant, we want to know the probability that the true outcome should be complete, that is, $P(Y_{i}=1)$. Let $\boldsymbol{x}_{i}$ be the set of original features of the $i^{th}$ applicant’s and $x_{i}^{\ast}$ be the assigned (synthetic) gender or university. Note, we drop the subscript $i$ when referring to a general applicant. We can write the probability of the true outcome for applicant with features $(\boldsymbol{x},x^{\ast})$ as $P(Y=1\mid\boldsymbol{x},x^{\ast})$ $=\frac{P(x^{\ast}\mid Y=1,\boldsymbol{x})P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$ Since we assign the values of the protected or proxy feature $x^{\ast}$ based solely on the ground-truth outcome, we can assume that $x^{\ast}$ and $\boldsymbol{x}$ are independent given $y$. So we say that $P(x^{\ast}\mid Y=1,\boldsymbol{x})=P(x^{\ast}\mid Y=1)$. $P(Y=1\mid\boldsymbol{x},x^{\ast})$ $=\frac{P(x^{\ast}\mid Y=1)P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$ $=\frac{P(Y=1\mid x^{\ast})P(x^{\ast})}{P(Y=1)}\frac{P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$ Then, removing any terms not containing $Y$, which can be normalized away, we are left with $P(Y=1\mid\boldsymbol{x},x^{\ast})$ $=\frac{P(Y=1\mid x^{\ast})P(Y=1\mid\boldsymbol{x})}{P(Y=1)}$ In the protected case, we know the probability of acceptance given the synthetic feature (that is, $P(Y=1\mid x^{\ast})$) based on our selected $60/40$ male/female acceptance ratio. In the proxy case, we estimate the probability of acceptance given the university using the joint distribution of gender and university and the probability of acceptance given gender. For each applicant, we estimate the probability of the applicant completing their loan $P(Y=1\mid\boldsymbol{x})$ based on only the non-synthetic features ($\boldsymbol{x}$) using a linear regression model that does not have access to the synthetic information. Finally, we can calculate $P(Y=1)$ based on the rate of ground-truth acceptances in the original data.

### v2

@@ 7.1. Justification Involving Protected Feature
Here, we analyze justifications that explicitly mention gender in the direct bias conditions. In these justifications, we assess how participants incorporate gender into their judgment of AI predictions or when making their own predictions. Pre-disclosure, when explanations are not provided, participants rarely discuss gender as a salient part of their justification. However, when explanations are provided, participants often mention trying to “ignore gender” when making their decision. Notably, participants who mentioned gender in predictions about female applicants tended to make “Complete” or “Neutral” predictions. Thus, even though explanations significantly decreased gender parity (§6.1) overall, they appear to help participants correct model biases in some cases.
Post-bias disclosure, justifications mentioning gender still predominantly appear in the condition with explanations. Nevertheless, some participants, even without explanations, mention gender bias and flip “Late” predictions for female applicants. For instance, one participant explained overriding such a prediction based on observing a male applicant with the same occupation predicted as “Complete”. In the condition with explanations, many participants asserted that “gender should not be a deciding factor” and ignored gender when making their prediction. Some participants, when supporting a ”Late” prediction for a female applicant, clarify that their decision was based on other features (“The large amount of negatives aside from gender still point towards being late.”). However, even in the case of direct bias with both explanations and bias disclosure, some participants still align with model biases. For example, one participant agreed with a “Complete” prediction of a male applicant “[b]ecause according to AI, male gender is more likely to complete loan…
@@ 7.2. Justifications Involving Proxy Feature
Here, we analyze justifications that explicitly mention universities in the indirect bias conditions and assess how participants incorporate universities in their judgments. Pre-disclosures, some participants mentioned that attending college generally increases the likelihood of repayment, regardless of the specific school (for example, mentioning that an applicant is “college-educated” and predicting “Complete”). However, other participants factor in the specific college when deciding to accept or reject an applicant. In conditions without explanations, we see evidence of participants relying on their own judgment of school quality. For instance, participants mention that they have “never heard of Kenyon College,” or that a co-ed school in the application “…is not a particularly prestigious university” (even though in the underlying model, attending a co-ed school is counted as a positive). In contrast, with explanations provided, we instead see examples of participants aligning their evaluation of a school with the model’s biases. For instance, on applicants from women’s colleges, participants claimed “The applicant didn’t go to a good college,” or “…College history was a major contributing factor to being late on loan.” On applications from co-ed schools, participants claimed that the applicant “…attended a good university,” or “the university is listed as a good one.” This supports our quantitative finding that explanations alone lead participants to align with model biases (§6.1).
After bias disclosure alone (and especially without explanations), mentions of universities were quite sparse. Some participants mentioned that the university feature is given excessive weight (“I just find it hilarious that the borrower’s state and university is such a huge factor.”) but may not recognize this as indirect bias. One participant, although aware that Bryn Mawr is a women’s college, expressed uncertainty about identifying biased predictions without direct access to protected features: “After learning more about possible discriminatory predictions on the AI’s part… I’m specifically concerned about gender and race… but don’t quite know how to discern that from these charts… This applicant profile gave me pause because I *think* Bryn Mawr College is an all-women’s college.
After full bias and correlation disclosure, the “university” feature appears frequently in justifications. Here, participants continue to highlight the excessive weight assigned to the university in explanations, noting particularly unwarranted negative weight on women’s colleges (e.g., “While the system said late, I thought this was unfair because it placed a strong negative value on the college, which might be a women’s college.”). Although some participants use this university bias as a justification for flipping model predictions, many acknowledge the bias and either make a neutral prediction or concur with predictions of women being late in repayment. Additionally, we observed that some participants struggled to recall which universities were co-ed, which may have limited their ability to intervene and correct model biases.
@@ 7.3. Justifications Indicating Over-reliance
In addition to the positive examples of explanations and disclosures helping participants notice and correct model biases, we also observe instances where decisions were solely based on the AI prediction or the corresponding explanations, regardless of disclosures. For example, even after bias disclosure in a direct bias condition, a participant agreed with a prediction of a female applicant being “Late” saying that “They seem to have more negatives than positives.” Similarly, after bias and correlation disclosures in an indirect bias condition, a participant changed the prediction for an applicant from a women’s college from “Complete” to “Late,” providing a similar justification. This indicates some participants persist in using explanations containing known biases, since for these applicants, ignoring the biased features (gender and university, respectively) would have resulted in the positives outweighing the negatives.
We also find instances of over-trust in AI even after being told that AI is biased such as “I have no reason to disagree with the AI, if the AI is discriminating it probably has a good reason to,” or “An AI is usually better than a professional let alone an amateur like me.” This indicates that even despite explanations and disclosures, there is room for improvement in educating and training humans to avoid unwarranted trust in AI systems and promote fair decision-making.
@@ 8. Discussion and Limitations
Explanations Only Disclosures Without Explanations Disclosures With Explanations Joint Intervention
Prot Prox Prot BD Prox BD Prox BD+CD Prot BD Prox BD Prox BD+CD Prot BD Prox BD Prox BD+CD
Fairness Rating $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$ $\downarrow$
Fairness Perception Fairness Saliency $\uparrow$ $\cdot$ $\cdot$ $\cdot$ $\uparrow$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\cdot$ $\uparrow$
DM Fairness Gender Parity $\downarrow$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\cdot$
Accuracy $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\downarrow$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$
FNR $\uparrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\downarrow$ $\cdot$ $\cdot$ $\cdot$ $\cdot$ $\cdot$
DM Quality FPR $\downarrow$ $\uparrow$ $\cdot$ $\cdot$ $\uparrow$ $\uparrow$ $\uparrow$ $\uparrow$ $\cdot$ $\cdot$ $\cdot$
Table 1. Summary of our main results. Arrows represent significant effects and point in the direction of the change. “BD” and “CD” represent bias and correlation disclosures, respectively.
In this work, we studied the effect of explanations and disclosures on fairness perceptions and decision-making when humans are provided predictions from models exhibiting direct or indirect bias. Our findings are summarized in Table 1. Regardless of intervention, we consistently observed that human-AI teams made fairer decisions than the AI alone. We found that explanations alone significantly improved participants’ ability to notice unfairness in the case of direct bias only. However, explanations led participants to be more influenced by model biases, whether they noticed these biases or not. Disclosures were an effective tool for helping users recognize unfairness in the case of indirect bias, especially with the help of explanations. And we saw that this increased recognition of bias was paired with fairer human-AI decisions, showing that disclosures helped participants understand when and how to intervene on model decisions to produce fairer outcomes.
However, we found that the joint intervention of including both explanations and disclosures (over including neither) was only effective in helping participants recognize model bias, not correct it. If the main objective is to help users notice direct model biases, we recommend including explanations, and if it is to help users notice indirect model bias, we recommend including explanations and disclosing both model bias and the correlations between protected and proxy features. But if the main objective is to help the human-AI team produce fairer outcomes, we did not find including explanations with disclosures to be an effective intervention. However, if explanations are to be used, then disclosures may help contextualize explanations and the potential biases, especially when these biases are indirect. While in a perfect world, such known biases could be addressed in the model itself instead of relying on human intervention, this may not always be possible. In many cases, we may have limited access to the underlying model (e.g., only having API access) or may not be able to non-superficially “debias” it (Gonen and Goldberg, 2019). Disclosures may help uncover these biases to humans, possibly leading to fairer human-AI decisions.
A key limitation of work is that since we show our participants partially-synthetic loan data, we cannot directly rely on the existing ground truth. Instead, we calculate the expectation of ground-truth based metrics (accuracy, FNR, and FPR) which means that there are applicants for which neither choice is very likely to be “correct” (i.e., both $P(Y_{i}=1)$ and $P(Y_{i}=0)$ are close to $0.5$). We handle this in part by adjusting for the baseline AI-only scores; however, using a fully non-synthetic dataset and original ground truth values may lead to cleaner results. This lack of a true ground-truth, in part, led us to use demographic parity which has been argued to be insufficient as a notion of fairness (Dwork et al., 2012).
There is potential concern about the use of a loan prediction task since the participants are not financial experts. As we discuss in §5, participants are shown a subset of the original Prosper features that we believe are relatively intuitive without more than a commonsense understanding of lending (e.g., size of the loan being requested and employment status). We also hope that a task mimicking the loan approval process is high-stakes enough to encourage more care from the crowd-workers in their decision-making. However, more work is needed to study how our findings generalize to settings with varied task stakes or domain expertise.
Another limitation is that our study design forces participants to make decisions one at a time, without seeing the entire pool of applicants. It is our hope that the percent/percentile information given for each feature gave participants a better sense of how each applicant’s profile compared to the general pool, even without seeing many profiles. However, we recognize that it may be difficult for participants to conceptualize what a “strong” or “weak” candidate looks like under this design. This may make it more difficult for participants who, for example, wish to increase the acceptance rate of women in phase 2 to decide which female applicants are “most deserving” of having their prediction flipped to “Complete”.
Despite these limitations, our work provides insights into the effect of explanations on fairness in human-AI decision-making, especially when the biases are indirect (through a proxy features). We conclude that neither explanations nor disclosures alone improve the fairness of decisions made by a human-AI team. Our findings serve to caution the wider community from treating explanations as a foolproof solution to human-AI collaborative decision-making: explanations may not always make model biases clear and may make people more prone to align with model biases, leading to less fair decisions. When people are repeatedly exposed to explanations that justify or rationalize biased predictions, they may begin to accept these biases as valid or even desirable, rather than critically questioning and challenging them. We highlight that explanations and disclosures in conjunction may be helpful to some extent. However, more work is needed to further examine how best to aid humans not only in identifying indirect model biases, but also in systematically correcting these biases.
@@ Acknowledgements.
We sincerely thank the CHI TRAIT and IUI reviewers, Md Naimul Hoque, and the members of the UMD CLIP and HCIL labs for their valuable feedback. Additionally, we would like to thank Solon Barocas, Jennifer Wortman Vaughan, Forough Poursabzi-Sangdeh, and Mahsan Nourani for their work on the version of the Prosper Loan dataset used in this paper.
@@ A.1. Probability of Loan Completion
In our study, each applicant $i$ is only evaluated by a single participant $j$ in a given condition, and each participant $j$ evaluates $20$ applicants across the two phases (§4.1). We represent the set of observed decisions as $S=\{(i,j)\mid\text{participant $j$ sees applicant $i$}\}$. For the $i^{\text{th}}$ applicant, we want to know the probability that the true outcome should be complete, that is, $P(Y_{i}=1)$. Let $\boldsymbol{x}_{i}$ be the set of original features of the $i^{th}$ applicant’s and $x_{i}^{\ast}$ be the assigned (synthetic) gender or university. Note, we drop the subscript $i$ when referring to a general applicant. We can write the probability of the true outcome for applicant with features $(\boldsymbol{x},x^{\ast})$ as
$P(Y=1\mid\boldsymbol{x},x^{\ast})=\frac{P(x^{\ast}\mid Y=1,\boldsymbol{x})P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$
Since we assign the values of the protected or proxy feature $x^{\ast}$ based solely on the ground-truth outcome, we can assume that $x^{\ast}$ and $\boldsymbol{x}$ are independent given $y$. Thus, $P(x^{\ast}\mid Y=1,\boldsymbol{x})=P(x^{\ast}\mid Y=1)$.
$P(Y=1\mid\boldsymbol{x},x^{\ast})$ $=\frac{P(x^{\ast}\mid Y=1)P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$
$=\frac{P(Y=1\mid x^{\ast})P(x^{\ast})}{P(Y=1)}\frac{P(Y=1\mid\boldsymbol{x})}{P(x^{\ast}\mid\boldsymbol{x})}$
Then, removing any terms not containing $Y$, which can be normalized away, we are left with
$P(Y=1\mid\boldsymbol{x},x^{\ast})=\frac{P(Y=1\mid x^{\ast})P(Y=1\mid\boldsymbol{x})}{P(Y=1)}$
In the protected case, we know the probability of acceptance given the synthetic feature (that is, $P(Y=1\mid x^{\ast})$) based on our selected $60/40$ male/female acceptance ratio. In the proxy case, we estimate the probability of acceptance given the university using the joint distribution of gender and university and the probability of acceptance given gender. For each applicant, we estimate the probability of the applicant completing their loan $P(Y=1\mid\boldsymbol{x})$ based on only the non-synthetic features ($\boldsymbol{x}$) using a linear regression model that does not have access to the synthetic information. Finally, we can calculate $P(Y=1)$ based on the rate of ground-truth acceptances in the original data.